Given: An 8-bit computer system with 4 MB of main memory. It has a 16K-words direct-mapped cache with a block size of 16 Bytemultiply hifhlighted number by 4 and please explain
Answer #1

Solution:

Given,

=>8-bit computer system

=>Size of main memory = 4 MB

=>Size of cache = 16 K words

=>Direct mapped cache

=>Block size = 16 bytes

=>Byte address = 5678 in decimal

=>Need to multiply byte address by 4.

Explanation:

Direct mapped cache:

Tag Cache index Block offset

8 bits                      10 bits                                      4 bits

Number of bits required for main memory address:

=>Number of bits required to express main memory address = log2(main memory size)

=>Number of bits required to express main memory address = log2(4 MB)

=>Number of bits required to express main memory address = log2(4*2^20 B) as 1 MB = 2^20 B

=>Number of bits required to express main memory address = log2(2^22 B)

=>Number of bits required to express main memory address = 22 bits

Number of bits required for cache index:

=>Size of cache = 16 K words

=>We know that system is 8-bits so 1 word = 8 bits

=>1 word = 1 byte as 1 byte = 8 bits

=>Size of cache = 16 KB

=>Number of blocks in the cache memory = Size of cache memory/size of cache block

=>Number of blocks in the cache memory = 16*1024 B/16 B as 1 KB = 1024 B

=>Number of blocks in the cache memory = 1024 blocks(0-1023)

=>Number of bits required for cache index = log2(number of blocks)

=>Number of bits required for cache index = log2(1024)

=>Number of bits required for cache index = log2(2^10)

=>Number of bits required for cache index = 10 bits

Number of bits required for block offset:

=>Number of bits required for block offset = log2(size of block)

=>Number of bits required for block offset = log2(16 B)

=>Number of bits required for block offset = log2(2^4 B)

=>Number of bits required for block offset = 4 bits

Number of bits required for tag:

=>Number of bits required for tag = 22 - 4 - 10

=>Number of bits required for tag = 8 bits

Calculating value of cache line number:

=>Mulitply the byte address by 4 = 4*5678

=>So calculated new address = 22712

=>In direct mapped cache byte address is mapped according to the relation: B mod P = i where B is the byte address of main memory, P is the number of blocks present in the cache memory and i the cache line number where the main memory block address is mapped.

=>22712 mod 1024 = 184 in decimal

=>Cache line number in hexadecimal = (B8)16

I have explained each and every part with the help of statements attached to it.